# Weyl metrizability of 3-dimensional projective structures and CR submanifolds

### v2

@@ 2.2. Conformal geometry in dimension 3
In this section we review the Cartan geometric description of conformal structures in dimension three. A conformal structure of signature $(p,q)$ on an $n$-dimensional manifold $M,$ where $n=p+q,$ is defined as the equivalence class of (pseudo-)Riemannian metrics of signature $(p,q)$ that are proportional via an everywhere positive function on $M,$ i.e. a sub-bundle $[g]\subset\mathrm{Sym}^{2}(T^{*}M)$ defined as
$[g]=\left\{\lambda^{2}g\ \vline\ \lambda\in C^{\infty}(M,\mathbb{R}\setminus\{0\})\right\},$
@@
Conformal structures of signature $(p,q)$ define Cartan geometries $(\mu\colon\mathcal{C}\to M,\psi)$ of type $(\mathrm{SO}(p+1,q+1),Q_{1})$ where $Q_{1}\subset\mathrm{SO}(p+1,q+1)$ is the parabolic subgroup that stabilizes a null line in $\mathbb{R}^{p+1,q+1}.$ With a choice of orientation, which will be important in this paper, one can define oriented conformal structures as Cartan geometries of type $(\mathrm{SO}^{o}(p+1,q+1),Q_{1})$ where $\mathrm{SO}^{o}(p+1,q+1)$ is the identity components of $\mathrm{SO}(p+1,q+1)$ and $Q_{1}$ is the stabilizes a null ray in $\mathbb{R}^{p+1,q+1}.$ As parabolic geometries they are among the so-called regular and normal parabolic geometries, which we will not elaborate on. The flat model for conformal structures is the $n$-dimensional hyperquadric $Q_{p,q}\subset\mathbb{P}^{n+1}$ defined by the standard inner product of signature $(p+1,q+1)$ on $\mathbb{R}^{p+1,q+1}.$
@@
(2.1) $\psi=\begin{pmatrix}-\theta^{0}&\xi_{2}&\xi_{1}&\xi_{0}&0\\ \omega^{2}&0&-\theta^{3}&-\varepsilon\theta^{2}&\varepsilon\xi_{2}\\ \omega^{1}&\theta^{3}&0&-\varepsilon\theta^{1}&\varepsilon\xi_{1}\\ \omega^{0}&\theta^{2}&\theta^{1}&0&\xi_{0}\\ 0&\varepsilon\omega^{2}&\varepsilon\omega^{1}&\omega^{0}&\theta^{0}\end{pmatrix}.$
As a result, the symmetric bilinear form
@@
is well-defined and the conformal structure on $M$ is given by $[s^{*}g]\subset\mathrm{Sym}^{2}(T^{*}M)$ for any section $s\colon M\to\mathcal{C}.$ When $\varepsilon$ is $1$ or $-1,$ the conformal structure is Riemannian or Lorentzian, respectively.
@@
Its Cartan curvature $\Psi\in\Omega^{2}(\mathcal{C},\mathfrak{so}(p+1,q+1))$ is
(2.3) $\Psi=\mathrm{d}\psi+\psi{\,{\wedge}\;}\psi=\begin{pmatrix}0&\Xi_{2}&\Xi_{1}&\Xi_{0}&0\\ 0&0&0&0&\varepsilon\Xi_{2}\\ 0&0&0&0&\varepsilon\Xi_{1}\\ 0&0&0&0&\Xi_{0}\\ 0&0&0&0&0\end{pmatrix},$
@@ 2.3. CR geometry in dimension 5
where $\mathcal{L}:\Lambda^{2}\mathscr{C}\to TN/\mathscr{C}$ is the Levi bracket of $\mathscr{C}$ defined at every $x\in N$ as
@@
Denoting the $\sqrt{\delta}$ and $-\sqrt{\delta}$ eigenspaces of $\mathcal{J}_{\delta}$ by ,¯$\mathscr{H},\overline{\mathscr{H}}\subset\mathbb{K}_{\delta}\mathscr{C}:=\mathscr{C}\otimes\mathbb{K}_{\delta},$ respectively, a $\delta$-CR structure is an almost $\delta$-CR structure for which $\mathscr{H}$ and $\overline{\mathscr{H}}$ are integrable.
@@
in which $\tilde{X}\in\Gamma(\mathscr{K})$, $\tilde{Y}\in\Gamma(\overline{\mathscr{H}})$, $\tilde{X}_{x}=X$, $\tilde{Y}_{x}=Y$ with its conjugate ¯$\overline{\mathcal{L}}_{2}:\overline{\mathscr{K}}\otimes\mathscr{H}\to\overline{\mathscr{H}}/\overline{\mathscr{K}}$ defined similarly. A $\delta$-CR structure is called 2-nondegenerate if $\mathcal{L}_{2}(X,\overline{\mathscr{H}})=0$ implies $X=0$ and similarly for the conjugate ¯$\overline{\mathcal{L}}_{2}$.
@@
Identifying $TN/\mathscr{C}$ at each point $x\in N$ with $\mathbb{R},$ the Levi bracket $\mathcal{L}$ when $\delta=-1$ corresponds to the imaginary part of a Hermitian inner product of signature $(p,q),p+q=n$ and when $\delta=1,$ after composition with the para-complex structure, gives a metric of signature $(n,n).$ The induced action of $\mathcal{J}_{\delta}$ on $\mathbb{K}_{\delta}\mathscr{C}:=\mathbb{K}_{\delta}\otimes\mathscr{C}$ results in the splitting $\mathbb{K}_{\delta}\mathscr{C}=\mathscr{H}\oplus\overline{\mathscr{H}},$ where $\mathscr{H}$ and $\overline{\mathscr{H}}$ are (para-)holomorphic and anti-(para-)holomorphic sub-bundles of $\mathbb{K}_{\delta}\mathscr{C}$ for $\mathcal{J}_{\delta},$ respectively. Note that as a result of the compatibility condition (2.6), in the para-CR case both $\mathscr{H}$ and $\overline{\mathscr{H}}$ are null with respect to the induced inner product on $\mathbb{K}_{1}\mathscr{C}\cong\mathscr{C}.$ Furthermore, the compatibility condition in the para-CR case implies that in the splitting $\mathscr{C}=\mathscr{H}\oplus\overline{\mathscr{H}}$ the integrable distributions $\mathscr{H}$ and $\overline{\mathscr{H}}$ are Lagrangian with respect to the induced symplectic structure on $\mathscr{C}.$ As a result, a para-CR structure is also referred to as an integrable contact Legendrian/Lagrangian/Lagrangean structure. In the definition above when $N$ is 3-dimensional the eigenspaces $\mathscr{H}$ and $\overline{\mathscr{H}}$ have rank one and, therefore, are automatically integrable. Furthermore, for 2-nondegenerate CR-structures ($\delta=-1$) the nondegeneracy condition on ¯$\overline{\mathcal{L}}_{2}$ follows automatically, unlike the para-CR case ($\delta=+1$). Lastly, when the Levi form is nondegenerate, the trace condition in (2.6) is automatic for $\delta=-1$, however when $\delta=1,$ it ensures that the eigenspaces $\mathscr{H},\overline{\mathscr{H}}$ have equal rank.
Solving the equivalence problem for nondegenerate CR structures goes back to [Car33, CM74, Tan76, Web78]. Here we recall such a solution in the case of 5-dimensional nondegenerate CR structures of signature (1,1) as regular and normal parabolic geometries.
@@
A 5-dimensional CR structure with Levi form of signature (1,1) is a Cartan geometry $(\mathcal{Q}\to N,\phi)$ of type $(\mathrm{SU}(2,2),P)$ where $P$ is the stabilizer of a null line in $\mathbb{C}^{4}.$ The Cartan connection and curvature can be expressed as
(2.8) $\gamma=\begin{pmatrix}-\phi^{0}_{0}&\pi&\sigma&\sigma_{0}\\ \beta&\phi^{1}_{1}&\mathrm{i}\phi^{1}_{2}&\overline{\sigma}\\ -\mathrm{i}\alpha&\mathrm{i}\phi^{2}_{1}&-\overline{\phi}^{1}_{1}&\overline{\pi}\\ \mathrm{i}\alpha^{0}&\mathrm{i}\overline{\alpha}&\overline{\beta}&\overline{\phi}^{0}_{0}\end{pmatrix},\quad\Gamma=\mathrm{d}\gamma+\gamma{\,{\wedge}\;}\gamma=\begin{pmatrix}-\Phi^{0}_{0}&\Pi&\Sigma&\mathrm{i}\Sigma_{0}\\ 0&\Phi^{1}_{1}&\mathrm{i}\Phi^{1}_{2}&\overline{\Sigma}\\ 0&\mathrm{i}\Phi^{2}_{1}&-\overline{\Phi}^{1}_{1}&\overline{\Pi}\\ 0&0&0&\overline{\Phi}^{0}_{0}\end{pmatrix}$
@@
modulo $\{\alpha^{0}\},$ for real-valued functions $C_{0},\cdots,C_{4}$ on $\mathcal{Q}.$ The fundamental (harmonic) curvature can be represented by a weighted binary holomorphic quartic $C\in\Gamma(\mathrm{Sym}^{4}\mathscr{C}^{*})\otimes(\bigwedge\nolimits^{2}\mathscr{H})^{-1}\otimes(\mathscr{C}^{\perp})^{\textstyle{\frac{1}{2}}},$ where, for any section $s\colon N\to\mathcal{Q},$
@@
The fundamental binary quartic of a indefinite CR 5-manifold can be interpreted as an obstruction for the integrability of an $\mathbb{S}^{1}$-bundle of Lagrangian 2-planes in $\mathscr{C}$ that are null with respect to the indefinite subconformal structure on $\mathscr{C},$ which was first observed in [Bry82]. More generally, almost CR structures of signature $(1,1)$ for which the binary quartic is zero are equipped with a 4-parameter family of foliations by $\alpha$-planes which is an analogue of half-flatness in dimension four and has been studied in-detail in [KM24].
More precisely, in an indefinite CR structure, in terms of the Cartan connection (2.8), the conformal symplectic structure $[\rho]\subset\bigwedge\nolimits^{2}\mathscr{C}^{*}$ and indefinite subconformal structure $[g]\subset\mathrm{Sym}^{2}(\mathscr{C}^{*})$ are given by
@@
where, by abuse of notation, we have suppressed the pull-back of the 1-forms in (2.8) by a section $s\colon N\to\mathcal{Q}.$ As a result, the set of 2-planes in $\mathscr{C}$ that are null with respect to $g$ define two $\mathbb{P}^{1}$-bundles $\mathcal{A},\mathcal{B}\subset\mathrm{Gr}(2,TN)$ over $N$ whose fibers at a point $x\in N$ can be parametrized as follows
(2.11) $\mathcal{A}_{x}:=$$\left\{\mathrm{Ker}\left\{\alpha^{0},a_{2}\alpha^{1}-a_{1}\beta^{1},a_{2}\alpha^{2}-a_{1}\beta^{2}\right\}\ |\ [a_{1}:a_{2}]\in\mathbb{R}\mathbb{P}^{1}\right\}\subset\mathrm{Gr}(2,T_{x}N),$
$\mathcal{B}_{x}:=$$\left\{\mathrm{Ker}\left\{\alpha^{0},b_{1}\alpha^{1}-b_{2}\alpha^{2},b_{1}\beta^{1}-b_{2}\beta^{2}\right\}\ |\ [b_{1}:b_{2}]\in\mathbb{R}\mathbb{P}^{1}\right\}\subset\mathrm{Gr}(2,T_{x}N).$
By analogy with 4-dimensional indefinite conformal structures, we refer to the family $\mathcal{A}_{x}$ and $\mathcal{B}_{x}$ as $\alpha$-planes and $\beta$-planes at $x.$ Furthermore, we note that all $\alpha$-planes are Lagrangian with respect to $[\rho].$ However, among $\beta$-planes there is no real Lagrangian plane. The complexified $\beta$-planes for which $[b_{1}:b_{2}]$ is $[1:\mathrm{i}]$ and $[1:-\mathrm{i}]$ correspond to ,¯$\mathscr{H},\overline{\mathscr{H}}\subset\mathscr{C},$ respectively, and are the only Lagrangian planes among the complexified $\beta$-planes, whose integrability encodes the complex structure on $\mathscr{C}$.
Evaluating the holomorphic quartic $C$ on the set of $\alpha$-planes at $x\in N$ results in a weighted binary quartic. More explicitly, if $P_{a}\in\mathcal{A}_{x}$ where $a=[a_{1}:a_{2}],$ then define
(2.12) $Q_{a}=C(P_{a},P_{a},P_{a},P_{a})=\lambda(C_{4}a_{1}^{4}+4C_{3}a_{1}^{3}a_{2}+6C_{2}a_{2}^{2}a_{1}^{2}+4C_{1}a_{1}a_{2}^{3}+C_{0}a_{2}^{4}),$
for some non-vanishing scalar function $\lambda$. The roots of the binary quartic $Q_{a},$ viewed as Lagrangian null planes at $x$, give rise to the following definition.
@@
Naturally, if the binary quartic $C$ is identically zero, i.e. the CR structure is flat, then every $\alpha$-plane is a principal null plane. If the binary quartic is non-zero, then at each point there are at most four principal null planes. These properties suggest a notion of Petrov type for 5-dimensional CR structures, based on the number of roots of the binary quartic and their multiplicity and reality. The only case that will be of importance in this article is the case when the quartic has a real root of multiplicity four, which would be the analogue of Petrov type $\mathsf{N}$ is the context of 4-dimensional Lorentzian conformal geometry.

### v3

@@ 2.2. Conformal geometry in dimension 3
In this section we review the Cartan geometric description of conformal structures in dimension three. A conformal structure of signature $(p,q)$ on an $n$-dimensional connected manifold $M,$ where $n=p+q,$ is defined as the equivalence class of (pseudo-)Riemannian metrics of signature $(p,q)$ that are proportional via an everywhere positive function on $M,$ i.e. a sub-bundle $g]\subset\mathrm{Sym}^{2}(T^{*}M)$ defined as
$g]=\left\{\lambda^{2}g\ \vline\ \lambda\in C^{\infty}(M,\mathbb{R}\setminus\{0\})\right\},$
@@
Conformal structures of signature $(p,q)$ define Cartan geometries $(\mu\colon\mathcal{C}\to M,\psi)$ of type $(\mathrm{SO}(p+1,q+1),Q_{1})$ where $Q_{1}\subset\mathrm{SO}(p+1,q+1)$ is the parabolic subgroup that stabilizes a null line in $\mathbb{R}^{p+1,q+1}.$ With a choice of orientation, which will be important in this paper, one can define oriented conformal structures as Cartan geometries of type $(\mathrm{SO}^{o}(p+1,q+1),Q_{1})$ where $\mathrm{SO}^{o}(p+1,q+1)$ is the identity components of $\mathrm{SO}(p+1,q+1)$ and $Q_{1}$ stabilizes a null ray in $\mathbb{R}^{p+1,q+1}.$ As parabolic geometries, they are among the so-called regular and normal parabolic geometries, which we will not elaborate on. The flat model for conformal structures is the $n$-dimensional hyperquadric $Q_{p,q}\subset\mathbb{P}^{n+1}$ defined by the standard inner product of signature $(p+1,q+1)$ on $\mathbb{R}^{p+1,q+1}.$
@@
(2.1) $\psi=\begin{pmatrix}-\theta^{0}&\xi_{2}&\xi_{1}&\xi_{0}&0\\ \omega^{2}&0&-\theta^{3}&-\varepsilon\theta^{2}&\varepsilon\xi_{2}\\ \omega^{1}&\theta^{3}&0&-\varepsilon\theta^{1}&\varepsilon\xi_{1}\\ \omega^{0}&\theta^{2}&\theta^{1}&0&\xi_{0}\\ 0&\varepsilon\omega^{2}&\varepsilon\omega^{1}&\omega^{0}&\theta^{0}\end{pmatrix}$,
where we are using the fact that the conformal geometry of metrics of signature $(p,q)$ and $(q,p)$ are equivalent since the Lie groups $\mathrm{SO}(p+1,q+1)$ and $\mathrm{SO}(q+1,p+1)$ are isomorphic. The symmetric bilinear form
@@
is well-defined and the conformal structure on $M$ is given by $s^{*}g]\subset\mathrm{Sym}^{2}(T^{*}M)$ for any section $s\colon M\to\mathcal{C}.$ When $\varepsilon$ is $1$ or 1,-1, the conformal structure is Riemannian or Lorentzian, respectively.
@@
Its Cartan curvature is
(2.3) $\mathrm{d}\psi+\psi{\,{\wedge}\;}\psi=\begin{pmatrix}0&\Xi_{2}&\Xi_{1}&\Xi_{0}&0\\ 0&0&0&0&\varepsilon\Xi_{2}\\ 0&0&0&0&\varepsilon\Xi_{1}\\ 0&0&0&0&\Xi_{0}\\ 0&0&0&0&0\end{pmatrix}$
@@ 2.3. CR geometry in dimension 5
where $\mathcal{L}:\bigwedge\nolimits^{2}\mathscr{C}\to TN/\mathscr{C}$ is the Levi bracket of $\mathscr{C}$ defined at every $x\in N$ as
@@
Denoting the $\sqrt{\delta}$ and $\sqrt{\delta}$ eigenspaces of $\mathcal{J}_{\delta}$ by $\mathscr{H},\overline{\mathscr{H}}\subset\mathbb{K}_{\delta}\mathscr{C}:=\mathscr{C}\otimes\mathbb{K}_{\delta},$ respectively, a $\delta$-CR structure is an almost $\delta$-CR structure for which $\mathscr{H}$ and $\overline{\mathscr{H}}$ are integrable.
@@
in which $\tilde{X}\in\Gamma(\mathscr{K})$, $\tilde{Y}\in\Gamma(\overline{\mathscr{H}})$, $\tilde{X}_{x}=X$, $\tilde{Y}_{x}=Y$ with its conjugate $\overline{\mathcal{L}}_{2}:\overline{\mathscr{K}}\otimes\mathscr{H}\to\overline{\mathscr{H}}/\overline{\mathscr{K}}$ defined similarly. A $\delta$-CR structure is called 2-nondegenerate if $\mathcal{L}_{2}(X,\overline{\mathscr{H}})=0$ implies $X=0$ and similarly for the conjugate $\overline{\mathcal{L}}_{2}$.
@@
Identifying $TN/\penalty 50\mathscr{C}$ at each point $x\in N$ with $\mathbb{R},$ the Levi bracket $\mathcal{L}$ when $\delta=-1$ corresponds to the imaginary part of a Hermitian inner product of signature $(p,q),p+q=n,$ on $\mathscr{C}\otimes\mathbb{C},$ defined up to a conformal factor. The real part of this sub-conformal Hermitian structure on $\mathscr{C}\otimes\mathbb{C}$ is given by $g(X,Y)=\mathcal{L}(X,\mathcal{J}_{-1}Y),$ for $X,Y\in\Gamma(\mathscr{C}),$ which also has signature $(p,q).$ When $\delta=1,$ after composition with the para-complex structure, $\mathcal{J}_{1}$, the symmetric bilinear form $g$ on $\mathscr{C}$ has signature $(n,n)$ and is the real part of a subconformal para-Hermitian structure. The induced action of $\mathcal{J}_{\delta}$ on $\mathbb{K}_{\delta}\mathscr{C}:=\mathbb{K}_{\delta}\otimes\mathscr{C}$ results in the splitting $\mathbb{K}_{\delta}\mathscr{C}=\mathscr{H}\oplus\overline{\mathscr{H}},$ where $\mathscr{H}$ and $\overline{\mathscr{H}}$ are (para-)holomorphic and anti-(para-)holomorphic sub-bundles of $\mathbb{K}_{\delta}\mathscr{C}$ for $\mathcal{J}_{\delta},$ respectively. Note that as a result of the compatibility condition (2.6), in the para-CR case both $\mathscr{H}$ and $\overline{\mathscr{H}}$ are null with respect to the induced inner product on $\mathbb{K}_{1}\mathscr{C}\cong\mathscr{C}.$ Furthermore, the compatibility condition in the para-CR case implies that in the splitting $\mathscr{C}=\mathscr{H}\oplus\overline{\mathscr{H}}$ the integrable distributions $\mathscr{H}$ and $\overline{\mathscr{H}}$ are Lagrangian with respect to the induced symplectic structure on $\mathscr{C}.$ As a result, a para-CR structure is also referred to as an integrable contact Legendrian/Lagrangian/Lagrangean structure. In the definition above when $N$ is 3-dimensional the eigenspaces $\mathscr{H}$ and $\overline{\mathscr{H}}$ have rank one and, therefore, are automatically integrable. Furthermore, for 2-nondegenerate CR-structures ($\delta=-1$) the nondegeneracy condition on $\overline{\mathcal{L}}_{2}$ follows automatically, unlike the para-CR case ($\delta=+1$). Lastly, when the subconformal $\delta$-Hermitian structure on $\mathscr{C}$ is nondegenerate, the trace condition in (2.6) is automatic for $\delta=-1$, however when $\delta=1,$ it ensures that the eigenspaces $\mathscr{H},\overline{\mathscr{H}}$ have equal rank.
Solving the equivalence problem for nondegenerate CR structures goes back to [Car33, CM74, Tan76, Web78]. Here we recall such a solution in the case of 5-dimensional nondegenerate CR structures of signature (1,1) as regular and normal parabolic geometries, which is a special case of what appears in [CM74, Equation (5.33) and Appendix] wherein the term Cartan geometry does not appear.
@@
A 5-dimensional CR structure with Levi form of signature (1,1) is a Cartan geometry $(\mathcal{Q}\to N,\phi)$ of type $(\mathrm{SU}(2,2),P)$ where $P$ is the stabilizer of a null line in $\mathbb{C}^{2,2}.$ The Cartan connection and curvature can be expressed as
(2.8) $\gamma=\begin{pmatrix}-\phi^{0}_{0}&\pi&\sigma&\sigma_{0}\\ \beta&\phi^{1}_{1}&\mathrm{i}\phi^{1}_{2}&\overline{\sigma}\\ -\mathrm{i}\alpha&\mathrm{i}\phi^{2}_{1}&-\overline{\phi}^{1}_{1}&\overline{\pi}\\ \mathrm{i}\alpha^{0}&\mathrm{i}\overline{\alpha}&\overline{\beta}&\overline{\phi}^{0}_{0}\end{pmatrix},\quad\mathrm{d}\gamma+\gamma{\,{\wedge}\;}\gamma=\begin{pmatrix}-\Phi^{0}_{0}&\Pi&\Sigma&\mathrm{i}\Sigma_{0}\\ 0&\Phi^{1}_{1}&\mathrm{i}\Phi^{1}_{2}&\overline{\Sigma}\\ 0&\mathrm{i}\Phi^{2}_{1}&-\overline{\Phi}^{1}_{1}&\overline{\Pi}\\ 0&0&0&\overline{\Phi}^{0}_{0}\end{pmatrix}$
@@
modulo $\{\alpha^{0}\},$ for real-valued functions $C_{0},\cdots,C_{4}$ on $\mathcal{Q}.$ The fundamental (harmonic) curvature, also referred to as the Chern-Moser curvature, can be represented by a weighted binary holomorphic quartic $C\in\Gamma(\mathrm{Sym}^{4}\mathscr{C}^{*})\otimes(\bigwedge\nolimits^{2}\mathscr{H})^{-1}\otimes(\mathscr{C}^{\perp})^{\textstyle{\frac{1}{2}}},$ where, for any section $s\colon N\to\mathcal{Q},$
@@
The fundamental binary quartic of a indefinite CR 5-manifold can be interpreted as an obstruction to the integrability of an $\mathbb{S}^{1}$-bundle of Lagrangian 2-planes in $\mathscr{C}$ that are null with respect to the indefinite subconformal structure on $\mathscr{C},$ as was first observed in [Bry82]. More generally, almost CR structures of signature $(1,1)$ for which the binary quartic above is zero has been studied in-detail in [KM24], where it is shown that they are equipped with a 4-parameter family of foliations by null and Lagrangian 2-planes, providing a 5-dimensional analogue of half-flatness in 4-dimensional conformal geometry,
Recall that in a non-degenerate CR structure of hypersurface type if $\alpha^{0}$ is a contact 1-form, i.e. $\mathscr{C}=\mathrm{Ker}\{\alpha^{0}\},$ then $\rho:=\mathrm{d}\alpha^{0}$ defines a symplectic 2-form on $\mathscr{C},$ which is related to the Levi-bracket via $\alpha^{0}(\mathcal{L}(X,Y))=\alpha^{0}([X,Y])=-\rho(X,Y).$ The corresponding sub-Hermitian metric, also referred to as the Levi form, is defined as $h(X,Y)=\frac{1}{2\mathrm{i}}\rho(X,\bar{Y})$ for $X,Y\in\Gamma(\mathscr{C}\otimes\mathbb{C}).$ Using the almost complex structure $\mathcal{J}\in\mathrm{Aut}(\mathscr{C}),$ with $\mathcal{J}^{2}=-\mathrm{Id},$ the symmetric bilinear form $g(X,Y):=\rho(X,\mathcal{J}Y),$ for $X,Y\in\Gamma(\mathscr{C}),$ is the real part of $h$ and has signature $(p,q).$ In general, since there is no distinguished contact 1-form, the symplectic 2-form, $\rho,$ and the symmetric bilinear form, $g,$ are defined up to a conformal factor. In CR structures with Levi form of signature (1,1), in terms of the Cartan connection (2.8), the conformal symplectic structure $\rho]\subset\bigwedge\nolimits^{2}\mathscr{C}^{*}$ and indefinite subconformal structure $g]\subset\mathrm{Sym}^{2}(\mathscr{C}^{*})$ are given by
@@
where, by abuse of notation, we have suppressed the pull-back of the 1-forms in (2.8) by a section $s\colon N\to\mathcal{Q},$ and, as stated in Theorem 2.5, $\alpha=\alpha^{1}+\mathrm{i}\alpha^{2},$ $\beta=\beta^{1}+\mathrm{i}\beta^{2}$ belong to the dual $\mathrm{i}$-eigenspace of $\mathcal{J},$ and are referred to as the holomorphic 1-form on $\mathscr{C}\otimes\mathbb{C}.$
As a result, the set of 2-planes in $\mathscr{C}$ that are null with respect to $g$ define two $\mathbb{P}^{1}$-bundles $\mathcal{A},\mathcal{B}\subset\mathrm{Gr}(2,TN)$ over $N$ whose fibers at a point $x\in N$ can be parametrized as follows
(2.11) $\mathcal{A}_{x}$= $\left\{\mathrm{Ker}\left\{\alpha^{0},a_{2}\alpha^{1}-a_{1}\beta^{1},a_{2}\alpha^{2}-a_{1}\beta^{2}\right\}\ |\ [a_{1}:a_{2}]\in\mathbb{R}\mathbb{P}^{1}\right\}\subset\mathrm{Gr}(2,T_{x}N),$
$\mathcal{B}_{x}$= $\left\{\mathrm{Ker}\left\{\alpha^{0},b_{1}\alpha^{1}-b_{2}\alpha^{2},b_{1}\beta^{1}-b_{2}\beta^{2}\right\}\ |\ [b_{1}:b_{2}]\in\mathbb{R}\mathbb{P}^{1}\right\}\subset\mathrm{Gr}(2,T_{x}N).$
By analogy with 4-dimensional indefinite conformal structures, we refer to the family $\mathcal{A}_{x}$ and $\mathcal{B}_{x}$ as $\alpha$-planes and $\beta$-planes at $x.$ Furthermore, we note that all $\alpha$-planes are Lagrangian with respect to $\rho].$ However, among $\beta$-planes there is no real Lagrangian plane. The complexified $\beta$-planes for which $b_{1}:b_{2}]$ is $1:\mathrm{i}]$ and $1:-\mathrm{i}]$ correspond to $\mathscr{H},\overline{\mathscr{H}}\subset\mathscr{C},$ respectively, and are the only Lagrangian planes among the complexified $\beta$-planes, whose integrability encodes the complex structure on $\mathscr{C}$.
Using the dual frame, every $\alpha$-plane $P_{a}\in\mathcal{A}_{x},$ $a=[a_{1}:a_{2}]$ is given by $\langle a_{1}\partial_{\alpha^{1}}+a_{2}\partial_{\beta^{1}},a_{1}\partial_{\alpha^{2}}+a_{2}\partial_{\beta^{2}}\rangle,$ and, thus, contains a unique contact holomorphic direction $\langle V_{a}\rangle\subset P_{a}^{\mathbb{C}}$ where
$V_{a}=a_{1}\partial_{\alpha}+a_{2}\partial_{\beta}\in\Gamma(\mathscr{H}_{x}),$
and $\partial_{\alpha}=\textstyle{\frac{1}{2}}(\partial_{\alpha^{1}}-\mathrm{i}\partial_{\alpha^{2}})$ and $\partial_{\beta}=\textstyle{\frac{1}{2}}(\partial_{\beta^{1}}-\mathrm{i}\partial_{\beta^{2}}).$ Evaluating the holomorphic quartic $C$ on holomorphic vector fields in complexified $\alpha$-planes at $x\in N$ results in a weighted binary quartic. More explicitly, if $a=[a_{1}:a_{2}]\in\mathbb{P}^{1},$ and $V_{a}\in P_{a}^{\mathbb{C}}$ is a contact holomorphic vector field, then one can define
(2.12) $Q_{a}=C(V_{a},V_{a},V_{a},V_{a})=\lambda(C_{4}a_{1}^{4}+4C_{3}a_{1}^{3}a_{2}+6C_{2}a_{2}^{2}a_{1}^{2}+4C_{1}a_{1}a_{2}^{3}+C_{0}a_{2}^{4}),$
for some non-zero scalar $\lambda$. By our discussion above, the roots of the binary quartic $Q_{a}$ determine at most four Lagrangian null planes at $x$, which are used in the following definition.
@@
Naturally, if the binary quartic $C$ is identically zero, i.e. the CR structure is flat, then every $\alpha$-plane is a principal null plane. If the binary quartic is non-zero, then at each point there are at most four principal null planes. These properties suggest a notion of Petrov type for 5-dimensional CR structures, based on the number of roots of the binary quartic and their multiplicity and reality. The only important case in this article is the case when the quartic has a real root of multiplicity four, which would be the analogue of Petrov type $\mathsf{N}$ is the context of 4-dimensional Lorentzian conformal geometry.